A truck is stopped at a stoplight. When the light turns green, it accelerates at 2.5 m/s2. At the same instant, a car passes the truck going 15 m/s. Where and when does the truck catch up with the car?


1
Expert's answer
2021-09-12T18:59:41-0400

Truck

"u_t=0 \\

v_t^2 = 2a_ts_t"

Car

"u_c= 15 \;m/s \\

S= u_ct_c \\

S = 15t_c"

Time taken by truck = Time taken by car

"t_t =t_c \\

v_t = at_t \\

S= \frac{1}{2}at^2 = \frac{1}{2}(2.5)t^2 \\

15t_k = \frac{1}{2}(2.5)t^2 \\

\frac{30}{2.5} = t_c \\

t_c=12\;s \\

S= 15 \times 12 = 180 \;m"

At time 12 sec and distance 180 m both vehicles will meet.